# Lee–Yang Zeros and Pseudocritical Drift in J-Q Néel–VBS Transitions

### 16 Mar 2026 v1

@@ Abstract
Introduction. Quantum magnets provide a controlled setting for studying quantum phase transitions beyond the standard Landau–Ginzburg–Wilson framework based on a single local order parameter [46, 10, 56, 55, 57]. A central example is the square-lattice transition between Néel antiferromagnet (AFM) and a columnar valence-bond solid (VBS), realized in sign-problem-free J-Q models and accessible to large-scale quantum Monte Carlo simulations [51, 40, 37, 53, 58, 42, 49, 62]. Within conventional Landau theory, a direct continuous transition between these phases is not generically expected without additional fine tuning, since the Néel and VBS phases break distinct symmetries [10, 46, 56, 55]. The deconfined quantum critical point (DQCP) scenario instead posits a continuous transition governed by fractionalized spinons coupled to an emergent gauge field, with the Néel order parameter arising as a spinon bilinear and the VBS order associated with monopole operators [56, 55, 34, 28, 57]. Complementary to continuum field-theory approaches, Lee–Yang theory characterizes phase transitions through the analytic structure of the partition function $Z$ in the complex plane, where thermodynamic singularities are tied to the accumulation of zeros [73, 33, 24, 6, 1, 63, 36, 23, 11, 65]. Importantly, Lee–Yang theory is naturally formulated at finite size and is therefore well suited to numerical simulations. For finite linear system size $L$ and inverse temperature $\beta$, Lee–Yang zeros are discrete and remain off the real axis, but their scaling with $L$ provides a sharp, nonperturbative diagnostic of transition order and critical behavior [17, 24, 26, 30, 3, 12, 65]. Lee–Yang zeros therefore provide a natural probe of the debated Néel–VBS criticality in J-Q models [53, 58, 42, 74, 39, 54, 27, 31, 62].
@@
Early quantum Monte Carlo studies of square-lattice $\mathrm{SU}(2)$ J-Q models found nearly size-independent crossings of Binder cumulants and correlation-length ratios, consistent with a putative continuous Néel–VBS transition [51, 40, 37, 53]. At the same time, conventional finite-size diagnostics in these models remain difficult to interpret [53, 41, 58, 39, 49, 62]. Crossing points can look almost size independent on moderate lattices, yet continue to drift as larger sizes are included, while exponents from standard finite-size fits depend noticeably on the fitting window, indicating that the asymptotic regime is not cleanly reached. Joint AFM–VBS order-parameter distributions also exhibit approximate emergent symmetry, which is suggestive but does not by itself distinguish true criticality from a long-lived pseudocritical regime [41, 42, 66, 57]. Conventional real-axis observables therefore provide important evidence, but do not directly isolate the finite-size singularity scale or decisively determine the character of the transition. Distinguishing a stable unitary fixed point from an anomalously weak first-order transition thus remains a central challenge for understanding Néel–VBS criticality in the J-Q family [28, 41, 58, 39, 57].
@@
In this Letter we study the Néel–VBS transition through Lee–Yang zeros computed within stochastic series expansion (SSE) quantum Monte Carlo [50, 61, 16, 52, 71, 72, 11, 65]. By evaluating complex generating functions in symmetry-resolved channels, we track the finite-size flow of the leading Lee–Yang zeros, which directly measures the singularity scale and distinguishes critical behavior from first-order coexistence [24, 44, 2, 7, 3, 4]. Using amplitude-free two-size estimators constructed from low-lying zeros [24, 26, 30, 12, 65], we analyze benchmark models together with the uniform J-Q3 and J-Q2 systems. While the benchmarks recover their expected behavior, J-Q2 and J-Q3 exhibit pronounced finite-size drift, indicating an extended pseudocritical regime. These results support a weakly first-order interpretation of the Néel–VBS transition and highlight the advantage of Lee–Yang zeros as a direct finite-size diagnostic of slow drift [41, 58, 21, 39, 57, 62].
@@
Methods and Models. For finite $L$ and inverse temperature $\beta$, a symmetry-preserving anchor Hamiltonian $H_{0}$ defines an ensemble that respects all microscopic symmetries, so phases and criticality are inferred from the $L$ dependence of symmetry-resolved fluctuations [32, 10, 46, 28]. We probe a channel specified by operator support (site, bond, or plaquette) and translation sector $\mathbf{Q}$, defined from a local density $O(x)$ at position $\mathbf{R}_{x}$ by the momentum projection $O_{\mathbf{Q}}\equiv\sum_{x}e^{\mathrm{i}\mathbf{Q}\cdot\mathbf{R}_{x}}O(x)$, which is a translation eigen-operator and does not mix with inequivalent $\mathbf{Q}$ sectors in correlators [64, 32, 10, 46]. We introduce an imaginary source $\mathrm{i}g_{\mathbf{Q}}$ coupled to $O_{\mathbf{Q}}$, $H(\mathrm{i}g_{\mathbf{Q}})=H_{0}-\mathrm{i}g_{\mathbf{Q}}O_{\mathbf{Q}}$, and work with the normalized generator $G_{L,\beta}(\mathrm{i}g_{\mathbf{Q}})\equiv Z_{L,\beta}(\mathrm{i}g_{\mathbf{Q}})/Z_{L,\beta}(0)$, whose finite-$L$ zeros encode the analytic structure of the complex $Z$ [73, 33, 18, 24, 1] and whose thermodynamic pinching yields free-energy nonanalyticities [73, 33, 18, 24]. We evaluate $G$ in SSE by reweighting configurations sampled from the real anchor ensemble [11, 13, 67, 14, 15, 68]: for each stored configuration $\mathcal{C}$ we form the ratio $R(\mathrm{i}g_{\mathbf{Q}};\mathcal{C})\equiv W(\mathcal{C};\mathrm{i}g_{\mathbf{Q}})/W(\mathcal{C};0)$, where $W$ is the SSE weight including the source-dependent diagonal matrix elements [50, 61, 52], and compute $G(\mathrm{i}g_{\mathbf{Q}})=\langle R(\mathrm{i}g_{\mathbf{Q}};\mathcal{C})\rangle_{0}$. Lee–Yang zeros are then the values of $\mathrm{i}g_{\mathbf{Q}}$ at which this estimator vanishes [11, 65].
@@
which hosts an $\mathrm{O}(3)$ quantum phase transition between a Néel AFM and a quantum-disordered dimer phase [69, 70, 38]. We probe the Néel channel at $\mathbf{Q}=(\pi,\pi)$ through $M^{z}=\sum_{\mathbf{r}}(-1)^{x+y}S^{z}_{\mathbf{r}}$, by adding the imaginary source term $\mathrm{i}hM^{z}$, with $h\in\mathbb{R}$. For each SSE configuration, we record the numbers of diagonal bond operators acting on the two antiparallel local spin configurations $(\uparrow_{i}\downarrow_{j})$ and $(\downarrow_{i}\uparrow_{j})$, denoted $n^{(s)}_{\uparrow\downarrow}$ and $n^{(s)}_{\downarrow\uparrow}$, respectively. [^1] More precisely, these are counts of diagonal SSE vertices in bond sector $s$ whose local bond state is $(\uparrow_{i}\downarrow_{j})$ or $(\downarrow_{i}\uparrow_{j})$. The bond endpoints are ordered as $i\in A$ and $j\in B$ on the bipartite square lattice, and $s=\mu\epsilon$ labels the bond orientation $\mu\in\{x,y\}$ and the two translation-related bond sublattices $\epsilon=\pm$ selected by the columnar staggering factor $(-1)^{r_{\mu}}$ [11, 69]. With real anchor denominators $J_{s}$ and $\alpha_{s}=2/(N_{c}J_{s})$ ($N_{c}=4$ on the square lattice) [48, 50], the per-configuration ratio factor is $R_{h}(\mathrm{i}h)=\prod_{s}\bigl(1-\mathrm{i}\alpha_{s}h\bigr)^{n^{(s)}_{\uparrow\downarrow}}\bigl(1+\mathrm{i}\alpha_{s}h\bigr)^{n^{(s)}_{\downarrow\uparrow}}.$
@@
where $P_{ij}\equiv\tfrac{1}{4}-\mathbf{S}_{i}\!\cdot\!\mathbf{S}_{j}$, and $\square^{\prime}$ denotes one of the two checkerboard plaquette subfamilies, for example the red shaded plaquettes in Fig. 1(b) [74, 59, 28]. This model exhibits a discontinuous transition between a Néel AFM and a plaquette-singlet-solid (PSS) phase [74, 59, 35]. Motivated by the emergent $\mathrm{O}(4)$ symmetry reported at the transition [74, 59], we probe Lee–Yang zeros for both the Néel and PSS order parameters. For the AFM channel, we couple an imaginary source $\mathrm{i}h$ to the staggered magnetization $M^{z}=\sum_{\mathbf{r}}(-1)^{x+y}S^{z}_{\mathbf{r}}$. To detect PSS order, we resolve plaquettes by checkerboard and row parities, $\epsilon_{\rm row}(\square^{\prime})=(-1)^{y(\square^{\prime})}$ [74, 59], and define $O^{\rm PSS}\equiv\sum_{\square^{\prime}}\epsilon_{\rm row}(\square^{\prime})\,\mathcal{O}(\square^{\prime}),$ where $\mathcal{O}(\square^{\prime})$ is the local plaquette density associated with the microscopic interaction. [^2] In the SSE implementation, $\mathcal{O}(\square^{\prime})$ corresponds to the local four-spin operator on an active checkerboard plaquette, namely the projector product $P_{ij}P_{kl}$ appearing in Eq. (2). The PSS estimator is therefore obtained by counting operator-string insertions of $Q$-vertices on the active checkerboard family with row parity $\epsilon_{\rm row}=\pm 1$. We denote these counts by $N_{Q+}$ and $N_{Q-}$, respectively. Equivalently, $N_{Q\pm}$ are the numbers of plaquette-operator insertions carrying the two signs selected by the row-staggering factor in $O^{\rm PSS}$. In the PSS channel, we introduce an imaginary source term $\mathrm{i}\eta\,O^{\rm PSS}$. For each SSE configuration, the corresponding reweighting factor is determined by the real anchor coupling $Q$ and by the row-parity-resolved numbers of $Q$-vertex insertions, $N_{Q+}$ and $N_{Q-}$, giving $R_{\mathrm{pss}}(\mathrm{i}\eta)=\left(1+\mathrm{i}\frac{\eta}{Q}\right)^{N_{Q+}}\left(1-\mathrm{i}\frac{\eta}{Q}\right)^{N_{Q-}}.$
@@
where $\langle ij,kl\rangle\in\square$ denotes the two parallel bonds of an elementary plaquette, in either orientation, and $\langle ij,kl,mn\rangle\in\square$ denotes three adjacent parallel bonds on a $2\times 3$ or $3\times 2$ rectangle [51, 37, 58]. Tuning $J/Q_{p}$, with $p\in\{2,3\}$, drives the transition between the Néel AFM and the columnar VBS [51, 40, 37, 53, 41, 58]. To detect VBS order, we probe the bond-supported columnar channels at $\mathbf{Q}=(\pi,0)$ and $(0,\pi)$, using $D_{\mu}=\sum_{\mathbf{r}}(-1)^{r_{\mu}}\,\mathbf{S}_{\mathbf{r}}\!\cdot\!\mathbf{S}_{\mathbf{r}+\hat{\mu}}$. Motivated by the emergent $\mathrm{SO}(5)$ structure discussed for these transitions [41, 42, 66, 58, 49, 62, 57], one may equivalently probe the AFM sector or use plaquette-based definitions of VBS order; here we focus on the bond VBS channel and summarize alternative probes in the Supplemental Material. Implementing the imaginary source term $\mathrm{i}\xi_{\mu}D_{\mu}$, we separate bond insertions of orientation $\mu$ into the two translation-related subclasses distinguished by the sign of $(-1)^{r_{\mu}}$, and denote the corresponding counts by $N_{J+}$ and $N_{J-}$ [11]. [^3] In the SSE representation, $N_{J+}$ and $N_{J-}$ are the counts of diagonal $J$-bond operator insertions on bonds of orientation $\mu$ belonging to the two subclasses with $(-1)^{r_{\mu}}=+1$ and $(-1)^{r_{\mu}}=-1$, respectively. The resulting factorized ratio contribution is [11] $R_{J\mu}(\mathrm{i}\xi_{\mu})=\left(1+\mathrm{i}\frac{\xi_{\mu}}{J}\right)^{N_{J+}}\left(1-\mathrm{i}\frac{\xi_{\mu}}{J}\right)^{N_{J-}}.$
@@
At a continuous transition in a given symmetry sector, we denote by $\mathrm{i}g_{\mathbf{Q}}$ the imaginary source coupled to the corresponding order parameter $O_{\mathbf{Q}}$ and $g_{\mathrm{LY}}^{(k)}(L)$ the $k$th Lee–Yang zero in the complex source plane for linear system size $L$. A more detailed discussion of this scaling is given in the Supplemental Material. Since $g_{\mathbf{Q}}$ is the scaling field conjugate to $O_{\mathbf{Q}}$, with renormalization-group eigenvalue $y_{h}$ [17, 46, 9, 24], finite-size scaling gives $g_{\mathrm{LY}}^{(k)}(L)\sim A_{k}\,L^{-y_{h}},$ up to nonuniversal metric factors and irrelevant-operator corrections [24, 29, 26, 30, 12], with nonuniversal amplitude $A_{k}$. Using $y_{h}=(d+z+2-\eta)/2,$ where $d$ is the spatial dimension, $z$ is the dynamical exponent, and $\eta$ is the anomalous dimension of the order parameter, the Lee–Yang zeros therefore provide direct access to $\eta$ [46, 9]. At phase coexistence, by contrast, a small source biases competing phases by an amount proportional to the Euclidean spacetime volume $V_{\tau}=\beta L^{d}$, where $\beta$ is the inverse temperature, since the source term is summed over all sites and imaginary time. The zeros then scale as $g_{\mathrm{LY}}^{(k)}(L,\beta)\sim B_{k}/(\beta L^{d}),$ with nonuniversal amplitude $B_{k}$ [44, 2, 7, 3, 4, 5]. To remove $B_{k}$, we use the two-size estimator [^4] Starting from $g_{\mathrm{LY}}^{(k)}(L)\sim A_{k}L^{-y_{h}}$, the ratio for a size pair $(L,2L)$ eliminates $A_{k}$, $g_{\mathrm{LY}}^{(k)}(2L)/g_{\mathrm{LY}}^{(k)}(L)\sim 2^{-y_{h}},$ so $y_{h}=-\ln[g_{\mathrm{LY}}^{(k)}(2L)/g_{\mathrm{LY}}^{(k)}(L)]/\ln 2.$ Combining this with $y_{h}=(d+z+2-\eta)/2$ gives the estimator below. At finite size, corrections to scaling make this a size-dependent effective estimate of $\eta$, rather than its asymptotic value. [38, 12, 65] $\eta_{\mathrm{eff}}^{(k)}(L)=d+z+2+\frac{2}{\ln 2}\ln\!\left[\frac{g_{\mathrm{LY}}^{(k)}(2L)}{g_{\mathrm{LY}}^{(k)}(L)}\right].$ For $d=2$ and $z=1$, $\eta_{\mathrm{eff}}^{(k)}(L)$ approaches $\eta$ at a continuous transition, whereas for the first-order benchmarks studied here, with $\beta\propto L$, the volume law $g_{\mathrm{LY}}^{(k)}\sim L^{-(d+1)}$ gives the diagnostic value $\eta=-1$ [24, 44, 2, 7, 3, 4]. If the probed channel is not the ordering channel, the free energy remains analytic at $g_{\mathbf{Q}}=0$ in the thermodynamic limit, and the Lee–Yang zeros terminate at a finite edge [19, 9, 20, 1, 45]. In J-Q models, this gives a finite VBS-source (Néel-source) edge in the Néel (VBS) phase, closing only at the transition. We therefore focus on leading-zero scaling near criticality and leave edge behavior and its finite-size crossover to the Supplemental Material [24, 29, 25, 1].
@@
This perspective also clarifies why the leading Lee–Yang zeros can reveal slow pseudocritical drift more cleanly than real-axis observables in a weakly first-order regime [21, 39]. At first order the correlation length remains finite, so real-axis correlators do not approach a unique critical power law at coexistence and are strongly affected by phase mixing and analytic backgrounds [44, 2, 7]. By contrast, the leading Lee–Yang zeros directly encode the finite-size singularity scale through their distance from the real axis. In the coexistence regime they scale as $g_{\mathrm{LY}}(L)\sim L^{-(d+z)}$, corresponding to $\eta\to-1$, or equivalently to a vanishing effective scaling dimension of the order-parameter field, $\Delta_{\phi}=({d+z-2+\eta})/{2}\to 0$ [3, 4, 5, 26]. Since real-axis response functions are derivatives of $\ln Z$, they are most sensitive when a zero lies nearby. The leading zeros $g_{\mathrm{LY}}(L)$ can therefore reveal the slow crossover toward spacetime-volume scaling on size windows where conventional diagnostics still appear quasi-critical [12, 65].

### 30 Mar 2026 v2

@@ Abstract
Introduction. Quantum magnets provide a controlled setting for studying quantum phase transitions beyond the standard Landau–Ginzburg–Wilson framework based on a single local order parameter [65, 12, 75, 74, 76]. A central example is the square-lattice transition between Néel antiferromagnet (AFM) and a columnar valence-bond solid (VBS), realized in sign-problem-free J-Q models and accessible to large-scale quantum Monte Carlo simulations [70, 57, 53, 72, 78, 60, 68, 84]. Within conventional Landau theory, a direct continuous transition between these phases is not generically expected without additional fine tuning, since the Néel and VBS phases break distinct symmetries [12, 65, 75, 74]. The deconfined quantum critical point (DQCP) scenario instead posits a continuous transition governed by fractionalized spinons coupled to an emergent gauge field, with the Néel order parameter arising as a spinon bilinear and the VBS order associated with monopole operators [75, 74, 43, 36, 76, 63, 77, 55, 83, 94, 15, 107, 14, 104, 47, 79, 86, 87, 13]. Complementary to continuum field-theory approaches, Lee–Yang theory characterizes phase transitions through the analytic structure of the partition function $Z$ in the complex plane, where thermodynamic singularities are tied to the accumulation of zeros [103, 42, 32, 4, 85, 49, 31, 16, 92, 45, 46, 39, 90, 89, 91, 29, 3, 1, 22, 100, 48, 27, 2]. Importantly, Lee–Yang theory is naturally formulated at finite size and is therefore well suited to numerical simulations. For finite linear system size $L$ and inverse temperature $\beta$, Lee–Yang zeros are discrete and remain off the real axis, but their scaling with $L$ provides a sharp, nonperturbative diagnostic of transition order and critical behavior [23, 32, 34, 38, 6, 17, 92, 51, 50, 52, 58]. Lee–Yang zeros therefore provide a natural probe of the debated Néel–VBS criticality in J-Q models [72, 78, 60, 105, 56, 73, 35, 40, 84].
@@
Early quantum Monte Carlo studies of square-lattice $\mathrm{SU}(2)$ J-Q models found nearly size-independent crossings of Binder cumulants and scaled correlation length, consistent with a putative continuous Néel–VBS transition [70, 57, 53, 72]. At the same time, conventional finite-size diagnostics in these models remain difficult to interpret [72, 59, 78, 56, 68, 84]. Crossing points can look almost size independent on moderate lattices, yet continue to drift as larger sizes are included, while exponents from standard finite-size fits depend noticeably on the fitting window, indicating that the asymptotic regime is not cleanly reached. Joint AFM–VBS order-parameter distributions also exhibit approximate emergent symmetry, which is suggestive but does not by itself distinguish true criticality from a long-lived pseudocritical regime [59, 60, 93, 76]. Conventional real-axis observables therefore provide important evidence, but do not directly isolate the finite-size singularity scale or decisively determine the character of the transition. Distinguishing a stable unitary fixed point from an anomalously weak first-order transition thus remains a central challenge for understanding Néel–VBS criticality in the J-Q family [36, 59, 78, 56, 76, 96, 108].
@@
In this Letter we study the Néel–VBS transition through Lee–Yang zeros computed within stochastic series expansion (SSE) quantum Monte Carlo [69, 82, 21, 71, 101, 102, 16, 92]. By evaluating complex generating functions in symmetry-resolved channels, we track the finite-size flow of the low-lying Lee–Yang zeros, which directly measures the singularity scale and distinguishes critical behavior from first-order coexistence [32, 62, 5, 9, 6, 7]. Using amplitude-free two-size estimators constructed from low-lying zeros [32, 34, 38, 17, 92], we analyze benchmark models together with the uniform J-Q3 and J-Q2 systems. While the benchmarks recover their expected behavior, J-Q2 and J-Q3 exhibit pronounced finite-size drift, indicating an extended pseudocritical regime. These results support a weakly first-order interpretation of the Néel–VBS transition and highlight the advantage of Lee–Yang zeros as a direct finite-size diagnostic of slow drift [59, 78, 28, 56, 76, 84].
@@
Methods and Models. For finite $L$ and inverse temperature $\beta$, a symmetry-preserving anchor Hamiltonian $H_{0}$ defines an ensemble that respects all microscopic symmetries, so phases and criticality are inferred from the $L$ dependence of symmetry-resolved fluctuations [41, 12, 65, 36]. We probe a channel specified by operator support (site, bond, or plaquette) and translation sector $\mathbf{Q}$, defined from a local density $O(x)$ at position $\mathbf{R}_{x}$ by the momentum projection $O_{\mathbf{Q}}\equiv\sum_{x}e^{\mathrm{i}\mathbf{Q}\cdot\mathbf{R}_{x}}O(x)$, which is a translation eigen-operator and does not mix with inequivalent $\mathbf{Q}$ sectors in correlators [88, 41, 12, 65]. We introduce an imaginary source $\mathrm{i}g_{\mathbf{Q}}$ coupled to $O_{\mathbf{Q}}$, $H(\mathrm{i}g_{\mathbf{Q}})=H_{0}-\mathrm{i}g_{\mathbf{Q}}O_{\mathbf{Q}}$, and work with the normalized generator $G_{L,\beta}(\mathrm{i}g_{\mathbf{Q}})\equiv Z_{L,\beta}(\mathrm{i}g_{\mathbf{Q}})/Z_{L,\beta}(0)$, whose finite-$L$ zeros encode the analytic structure of the complex $Z$ [103, 42, 24, 32, 4] and whose thermodynamic pinching yields free-energy nonanalyticities [103, 42, 24, 32]. [^1] For the formal scaling discussion we adopt the convention $H(\mathrm{i}g_{\mathbf{Q}})=H_{0}-\mathrm{i}g_{\mathbf{Q}}O_{\mathbf{Q}}$. In the explicit SSE implementation the effective sign of the imaginary deformation is channel dependent. These differences amount only to channel-dependent relabelings of the scan coordinate ($g\to-g$ where needed) and therefore do not affect the Lee–Yang zero set or its finite-size scaling. We evaluate $G$ in SSE by reweighting configurations sampled from the real anchor ensemble [16, 18, 95, 19, 20, 97]: for each stored configuration $\mathcal{C}$ we form the ratio $R(\mathrm{i}g_{\mathbf{Q}};\mathcal{C})\equiv W(\mathcal{C};\mathrm{i}g_{\mathbf{Q}})/W(\mathcal{C};0)$, where $W$ is the SSE weight including the source-dependent diagonal matrix elements [69, 82, 71], and compute $G(\mathrm{i}g_{\mathbf{Q}})=\langle R(\mathrm{i}g_{\mathbf{Q}};\mathcal{C})\rangle_{0}$. Lee–Yang zeros are then the values of $\mathrm{i}g_{\mathbf{Q}}$ at which this estimator vanishes [16, 92].
@@
which hosts an $\mathrm{O}(3)$ quantum phase transition between a Néel AFM and a quantum-disordered dimer phase [98, 99, 54]. We probe the Néel channel at $\mathbf{Q}=(\pi,\pi)$ through $M^{z}=\sum_{\mathbf{r}}(-1)^{x+y}S^{z}_{\mathbf{r}}$, by adding the imaginary source term $\mathrm{i}hM^{z}$, with $h\in\mathbb{R}$. For each SSE configuration, we record the numbers of diagonal bond operators acting on the two antiparallel local spin configurations $(\uparrow_{i}\downarrow_{j})$ and $(\downarrow_{i}\uparrow_{j})$, denoted $n^{(s)}_{\uparrow\downarrow}$ and $n^{(s)}_{\downarrow\uparrow}$, respectively. [^2] These counts correspond to diagonal SSE vertices in bond sector $s$, with local bond states $(\uparrow_{i}\downarrow_{j})$ or $(\downarrow_{i}\uparrow_{j})$. The bond endpoints $i\in A$ and $j\in B$ on the bipartite lattice are labeled by $s=\mu\epsilon$, where $\mu\in\{x,y\}$ is the bond orientation and $\epsilon=\pm$ denotes the translation-related sublattices, selected by the columnar staggering factor $(-1)^{r_{\mu}}$ [16, 98]. With real anchor denominators $J_{s}$ and $\alpha_{s}=2/(N_{c}J_{s})$ ($N_{c}=4$ on the square lattice) [67, 69], the per-configuration ratio factor is $R_{h}(\mathrm{i}h)=\prod_{s}\bigl(1-\mathrm{i}\alpha_{s}h\bigr)^{n^{(s)}_{\uparrow\downarrow}}\bigl(1+\mathrm{i}\alpha_{s}h\bigr)^{n^{(s)}_{\downarrow\uparrow}}.$
@@
where $P_{ij}\equiv\tfrac{1}{4}-\mathbf{S}_{i}\!\cdot\!\mathbf{S}_{j}$, and $\square^{\prime}$ denotes one of the two checkerboard plaquette subfamilies, for example the red shaded plaquettes in Fig. 1(b) [105, 80, 36]. This model exhibits a discontinuous transition between a Néel AFM and a plaquette-singlet-solid (PSS) phase [105, 80, 44]. Motivated by the emergent $\mathrm{O}(4)$ symmetry reported at the transition [105, 80], we probe Lee–Yang zeros for both the Néel and PSS order parameters. For the AFM channel, we couple an imaginary source $\mathrm{i}h$ to the staggered magnetization $M^{z}=\sum_{\mathbf{r}}(-1)^{x+y}S^{z}_{\mathbf{r}}$. To detect PSS order, we resolve plaquettes by checkerboard and row parities, $\epsilon_{\rm row}(\square^{\prime})=(-1)^{y(\square^{\prime})}$ [105, 80], and define $O^{\rm PSS}\equiv\sum_{\square^{\prime}}\epsilon_{\rm row}(\square^{\prime})\,\mathcal{O}(\square^{\prime}),$ where $\mathcal{O}(\square^{\prime})$ is the local plaquette density associated with the microscopic interaction. [^3] In the SSE implementation, $\mathcal{O}(\square^{\prime})$ corresponds to the four-spin operator $P_{ij}P_{kl}$ on an active checkerboard plaquette (Eq. (2)). The PSS estimator is obtained by counting $Q$-vertex insertions with row parity $\epsilon_{\rm row}=\pm 1$, denoted $N_{Q+}$ and $N_{Q-}$, which correspond to plaquette-operator insertions selected by the row-staggering factor. In the PSS channel, we introduce an imaginary source term $\mathrm{i}\eta\,O^{\rm PSS}$. For each SSE configuration, the corresponding reweighting factor is determined by the real anchor coupling $Q$ and by the row-parity-resolved numbers of $Q$-vertex insertions, $N_{Q+}$ and $N_{Q-}$, giving $R_{\mathrm{pss}}(\mathrm{i}\eta)=\left(1+\mathrm{i}\frac{\eta}{Q}\right)^{N_{Q+}}\left(1-\mathrm{i}\frac{\eta}{Q}\right)^{N_{Q-}}.$
@@
where $\langle ij,kl\rangle\in\square$ denotes the two parallel bonds of an elementary plaquette, in either orientation, and $\langle ij,kl,mn\rangle\in\square$ denotes three adjacent parallel bonds on a $2\times 3$ or $3\times 2$ rectangle [70, 53, 78]. Tuning $J/Q_{p}$, with $p\in\{2,3\}$, drives the transition between the Néel AFM and the columnar VBS [70, 57, 53, 72, 59, 78]. To detect VBS order, we probe the bond-supported columnar channels at $\mathbf{Q}=(\pi,0)$ and $(0,\pi)$, using $D_{\mu}=\sum_{\mathbf{r}}(-1)^{r_{\mu}}\,\mathbf{S}_{\mathbf{r}}\!\cdot\!\mathbf{S}_{\mathbf{r}+\hat{\mu}}$. Motivated by the emergent $\mathrm{SO}(5)$ structure discussed for these transitions [59, 60, 93, 78, 68, 84, 76], one may equivalently probe the AFM sector or use plaquette-based definitions of VBS order; here we focus on the bond VBS channel and summarize alternative probes in the Supplemental Material. Implementing the imaginary source term $\mathrm{i}\xi_{\mu}D_{\mu}$, we separate bond insertions of orientation $\mu$ into the two translation-related subclasses distinguished by the sign of $(-1)^{r_{\mu}}$, and denote the corresponding counts by $N_{J+}$ and $N_{J-}$ [16]. [^4] In the SSE representation, $N_{J+}$ and $N_{J-}$ count diagonal $J$-bond operator insertions on bonds of orientation $\mu$, corresponding to the two subclasses with $(-1)^{r_{\mu}}=\pm 1$. The resulting factorized ratio contribution is [16] $R_{J\mu}(\mathrm{i}\xi_{\mu})=\left(1+\mathrm{i}\frac{\xi_{\mu}}{J}\right)^{N_{J+}}\left(1-\mathrm{i}\frac{\xi_{\mu}}{J}\right)^{N_{J-}}.$
@@
At a continuous transition in a given symmetry sector, we denote by $\mathrm{i}g_{\mathbf{Q}}$ the imaginary source coupled to the corresponding order parameter $O_{\mathbf{Q}}$ and $g_{\mathrm{LY}}^{(k)}(L)$ the $k$th Lee–Yang zero in the complex source plane for linear system size $L$. A more detailed discussion of this scaling is given in the Supplemental Material. Since $g_{\mathbf{Q}}$ is the scaling field conjugate to $O_{\mathbf{Q}}$, with renormalization-group eigenvalue $y_{h}$ [23, 65, 11, 32], finite-size scaling gives $g_{\mathrm{LY}}^{(k)}(L)\sim A_{k}\,L^{-y_{h}},$ up to nonuniversal metric factors and irrelevant-operator corrections [32, 37, 34, 38, 17], with nonuniversal amplitude $A_{k}$. Using $y_{h}=(d+z+2-\eta)/2,$ where $d$ is the spatial dimension, $z$ is the dynamical exponent, and $\eta$ is the anomalous dimension of the order parameter, the Lee–Yang zeros therefore provide direct access to $\eta$ [65, 11]. [^5] Throughout this work, $\eta$ extracted from $y_{h}=(d+z+2-\eta)/2$ is used as a finite-size diagnostic. Only in a genuine critical regime does it coincide with the true anomalous dimension; in the coexistence regime, $g_{\mathrm{LY}}(L,\beta)\sim(\beta L^{d})^{-1}$ corresponds to the diagnostic limit $\eta\to-1$. At phase coexistence, by contrast, a small source biases competing phases by an amount proportional to the Euclidean spacetime volume $V_{\tau}=\beta L^{d}$, where $\beta$ is the inverse temperature, since the source term is summed over all sites and imaginary time. The zeros then scale as $g_{\mathrm{LY}}^{(k)}(L,\beta)\sim B_{k}/(\beta L^{d}),$ with nonuniversal amplitude $B_{k}$ [62, 5, 9, 6, 7, 8]. To eliminate $B_{k}$, we use the two-size estimator [^6] Starting from $g_{\mathrm{LY}}^{(k)}(L)\sim A_{k}L^{-y_{h}}$, the ratio for a size pair $(L,2L)$ eliminates $A_{k}$, $g_{\mathrm{LY}}^{(k)}(2L)/g_{\mathrm{LY}}^{(k)}(L)\sim 2^{-y_{h}},$ so $y_{h}=-\ln[g_{\mathrm{LY}}^{(k)}(2L)/g_{\mathrm{LY}}^{(k)}(L)]/\ln 2.$ Combining this with $y_{h}=(d+z+2-\eta)/2$ gives the estimator below. At finite size, corrections to scaling make this a size-dependent effective estimate of $\eta$, rather than its asymptotic value. [54, 17, 92] $\eta_{\mathrm{eff}}^{(k)}(L)=d+z+2+\frac{2}{\ln 2}\ln\left[\frac{g_{\mathrm{LY}}^{(k)}(2L)}{g_{\mathrm{LY}}^{(k)}(L)}\right].$ For $d=2$ and $z=1$, $\eta_{\mathrm{eff}}^{(k)}(L)$ approaches $\eta$ at a continuous transition, whereas for the first-order benchmarks studied here, with $\beta\propto L$, the volume law $g_{\mathrm{LY}}^{(k)}\sim L^{-(d+1)}$ gives the diagnostic value $\eta=-1$ [32, 62, 5, 9, 6, 7]. If the probed channel is not the ordering channel, the free energy remains analytic at $g_{\mathbf{Q}}=0$ in the thermodynamic limit, and the Lee–Yang zeros terminate at a finite edge [25, 11, 26, 4, 64]. In J-Q models, this gives a finite VBS-source (Néel-source) edge in the Néel (VBS) phase, closing only at the transition. We therefore focus on leading-zero scaling near criticality and leave edge behavior and its finite-size crossover to the Supplemental Material [32, 37, 33, 4].
@@
This perspective also clarifies why the leading Lee–Yang zeros can diagnose slow pseudocritical drift more cleanly than conventional real-axis observables in a weakly first-order regime [28, 56]. At first order, coexistence is governed by competition between phases rather than by a single scale-invariant state. Real-axis observables such as order-parameter moments or spin correlations therefore need not approach a unique critical power law and can be strongly affected by phase mixing and analytic backgrounds [62, 5, 9]. By contrast, the leading Lee–Yang zeros directly encode the finite-size singularity scale through their distance from the real axis, and in the coexistence regime they cross over to the characteristic spacetime-volume scaling of a first-order transition. [^7] Near coexistence, finite-volume observables are phase-weighted mixtures of competing states, while thermodynamic response functions, being derivatives of $\ln Z$, are controlled by the nearby Lee–Yang zeros. In the ordering channel this gives $g_{\mathrm{LY}}(L,\beta)\sim(\beta L^{d})^{-1}$. See Supplemental Material for details. The leading zeros can therefore reveal the slow crossover toward coexistence on size windows where conventional diagnostics still appear quasi-critical [17, 92].